OpenStudy (anonymous):

Check my answer please! I will give a medal! Simplify the expression. 3w - 10w. My answer: 3 - 10 = -7*w = -7w Final answer: -7w

OpenStudy (pixiedust1):

Correct!! Great job! :D

OpenStudy (anonymous):

I would agree with your answer. This is because 10w is being subtracted by 3w 3-10 = -3

OpenStudy (anonymous):

Thank you everyone!

OpenStudy (anonymous):

3w - 10w = (3-10)w=-7w is better written as your answer.

OpenStudy (pixiedust1):

No problem!! Cool profile pic btw! God Bless!